OpenStudy (anonymous):

The horizontal asymptotes of 7/x-2

OpenStudy (anonymous):

0

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

assuming that y=7/(x-2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} y = \frac{ \frac{ 7 }{ x } }{ 1-\frac{ 2 }{ x } }\]\[\lim_{x \rightarrow \infty} y=\frac{ 0 }{ 1 }\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} y=0\]